# Semantic Self-Distillation for Language Model Uncertainty

### v1

@@ 4 Future Work
Improving distillation fidelity. Our results indicate that SSD’s utility as a pre-generation hallucination signal is primarily limited by distillation fidelity: detection performance improves when the student accurately recovers the teacher’s prompt-induced uncertainty ordering. Promising directions include scaling the distillation dataset, enriching the student input beyond a single-token/single-layer representation, and distilling from stronger teachers (larger $S$, or higher-quality semantic targets). These extensions would clarify whether observed failures reflect an information bottleneck in the prompt representation or insufficient supervision for density estimation.
@@
Combining prior and posterior signals. SSD exposes complementary reliability signals from a single predicted density: pre-generation dispersion forecasts risk, while post-generation likelihood and distance-to-consensus verify prompt–answer agreement. A natural next step is to integrate these signals into decision policies for agents, for example abstaining or retrieving evidence under high dispersion, and accepting or rejecting candidate actions based on likelihood or consensus distance.
Diffusion models. While we focus on autoregressive models, SSD may be particularly well suited to diffusion-based language models, where hidden prompt representations may more directly parameterize the final semantic output distribution (Sahoo et al., 2024).
Domain-specific uncertainty estimation. The technique naturally extends to domain-specific foundation models, such as those in healthcare (Renc et al., 2024; Steinberg et al., 2024) and time-series forecasting (Faw et al., 2025). By distilling predictive distributions over latent outcome representations, students could provide real-time uncertainty estimates for future patient trajectories or other temporal processes, enabling risk-aware decision-making in sequential settings (Zeng et al., 2025).
@@ 5 Conclusion
We introduced Semantic Self-Distillation (SSD), a distributional probing framework that distils a language model’s sampled semantic answer distribution into a lightweight, prompt-conditioned density estimator. Unlike scalar probes which compress uncertainty into a single score, SSD predicts an explicit semantic density, enabling analytic dispersion scoring and additional distribution-level operations in a single forward pass.
Empirically, SSD provides competitive hallucination-risk signals relative to finite-sample baselines on several model families, while avoiding the inference-time overhead of generating multiple completions. Crucially, modelling the full density unlocks capabilities that scalar probes cannot match: the posterior likelihood provides strong context verification checks for out-of-domain answers, while mixture statistics allow for single-pass semantic consensus estimates.
Overall, SSD demonstrates that distilling a semantic predictive distribution can yield a general-purpose uncertainty object with applications in forecasting and verification. By decoupling these capabilities from expensive inference-time sampling, SSD offers a practical framework for integrating principled uncertainty quantification into latency-critical agentic applications.
6 Acknowledgments
@@ Appendix A Derivations
Appendix B Model Metadata
Table 5: Model characteristics. ‘Base Acc’ denotes QA accuracy. ‘PCP’ and ‘SEP’ indicate optimal probe layers. The ‘Best SSD Config’ columns show the student architecture (HH=Hidden, DD=Depth, KK=Components) that achieved the highest SSD AUROC.
Probe Layers Best SSD Config
Model Base Acc PCP SEP H D K AUROC
Qwen3-8B 0.620 32 24 128 6 10 0.677
Qwen3-4B-Instruct-2507 0.565 18 20 256 2 10 0.699
Llama 3.1 8B 0.700 20 16 256 4 2 0.687
Llama 3.2 3B 0.585 15 19 512 3 5 0.737
Ministral 8B 0.656 33 34 1024 6 1 0.748
SmolLM3 3B 0.567 20 19 512 6 5 0.710
Gemma 3 4B 0.532 22 19 512 2 10 0.700
Appendix C Computational Complexity Analysis
Sampling-based Methods (SE, TD).
Computing Semantic Entropy (Farquhar et al., 2024) requires $S$ stochastic sequences per prompt at inference time (typically $S\in[5,20]$). The cost is dominated by the autoregressive decoding steps:
where $\mathcal{O}_{\text{NLI}}$ represents the quadratic cost of comparing all sample pairs using an entailment model. This linear scaling with $S$ introduces latency that is often unacceptable for interactive agents.
Scalar Probes (PCP, SEP).
Semantic Self-Distillation (SSD).
Appendix D Ablations and Architecture Analysis
D.1 PCA Dimension
@@
This introduces a trade-off: higher dimensions preserve more semantic detail (increasing the theoretical upper bound of detection performance, measured by teacher dispersion), but make the regression task significantly harder for the student. As shown in Figure 3 and Table 6, while the detection performance of teacher dispersion improves monotonically with dimensionality, the student’s performance (SSD) peaks at $d_{pca}=16$. Beyond this point, the student fails to effectively distil the increasingly sparse density, leading to a drop in detection accuracy. For the OOD experiments we find PCA dimensions of 16 and 32 perform similarly, followed by a drop-off at higher dimensions. We fix $d_{pca}=16$ for all main experiments bar the OOD task, where we set $d_{pca}=32$.
@@
Table 6: Effect of PCA dimensionality on Hallucination Detection (SSD) and OOD Verification. Scores are averaged across the best student configuration for each model.
@@
Figure 3: As PCA dimension increases, the semantic resolution improves and the teacher dispersion baseline rises. The distillation task however becomes harder, causing SSD performance to fall.
D.2 Student Capacity and Information Saturation
@@
Table 5 reveals that the optimal hyperparameters vary by base model. Aggregating results across all models (Table 7) reveals capacity saturation, where increasing the network width or depth yields negligible gains in mean AUROC. This saturation suggests that the bottleneck is not the student’s expressivity, but the mutual information available in the final-token prompt representation. The student appears to have extracted most of the readily accessible signal; further improvements would require richer inputs (e.g., attention-pooled representations) rather than larger students. Although we observe optimal performance for $K=10$ mixture components, we opt not to increase this further given our set number of $S=32$ samples; modelling more than ten unique semantic concepts in this setting is unrealistic and would likely damage the generalization ability of the trained student.
Table 7: Marginal impact of student capacity (Fixed PCA=16). Scores represent the mean AUROC across all runs with the specified parameter; standard deviations indicate variability across different base models and remaining hyperparameter configurations. We observe negligible performance gains from increasing complexity, indicating information saturation in the prompt representation.
@@
Figure 4: Student capacity ablation, averaged across all seven models investigated.
Appendix E Full Hallucination Detection Results
Table 8: Hallucination Detection AUROC. Values report mean AUROC (%) over 1,000 bootstrap resamples of the test set, with the bootstrap standard deviation shown as a subscript. We mark the best overall result in bold and the best single-pass result with an underline. SSD achieves competitive performance with teacher dispersion but lags scalar probes and SE.
Model PCP SEP SSD (Ours) TD SE
Qwen3-8B 76.91.5 76.21.5 67.71.8 68.41.7 81.61.4
Qwen3-4B 80.21.4 78.31.5 69.91.6 73.21.6 78.51.5
Llama 3.1 8B 77.01.6 74.51.7 68.62.0 69.51.7 85.21.3
Llama 3.2 3B 76.71.5 75.71.5 73.81.7 66.51.7 80.71.4
Ministral 8B 74.01.7 77.11.5 74.81.6 74.31.6 80.51.5
SmolLM3 3B 79.71.4 78.91.4 71.11.6 69.81.6 83.21.2
Gemma 3 4B 75.41.5 70.71.7 70.11.7 67.61.7 73.11.6
Table 9: Hallucination Detection AUPRC. Values report mean AUPRC (%) over 1,000 bootstrap resamples of the test set, with the bootstrap standard deviation shown as a subscript. We mark the best overall result in bold and the best single-pass result with an underline. SSD consistently outperforms the teacher dispersion (TD) baseline, effectively shifting the cost of dispersion estimation from inference time to training time.
Model PCP SEP SSD (Ours) TD SE
Qwen3-8B 64.92.7 64.92.7 57.02.8 52.52.7 75.02.1
Qwen3-4B 74.32.3 70.82.5 61.82.5 59.62.5 73.32.1
Llama 3.1 8B 59.63.1 54.03.0 51.93.0 44.32.8 73.52.7
Llama 3.2 3B 68.92.4 65.42.6 65.72.7 55.82.7 74.22.4
Ministral 8B 59.22.9 62.92.8 59.92.9 57.92.9 69.62.7
SmolLM3 3B 75.12.2 73.32.2 62.32.6 56.92.5 78.91.9
Gemma 3 4B 69.42.4 65.02.5 66.22.5 63.22.4 69.22.3

### v2

@@ 5 Discussion & Future Work
Improving distillation fidelity. Our results in Appendix B.1 show a strong correlation between distillation fidelity and hallucination prediction performance, indicating that SSD’s effectiveness depends on how accurately the student reconstructs the teacher’s prompt-induced uncertainty ordering. Promising directions to improve this fidelity include scaling the distillation dataset, enriching the student input beyond a single-token/single-layer representation, and distilling from stronger teachers (for example sampling more responses per question, or using higher-quality semantic targets). Subsequent findings may clarify whether observed performance gaps between models stem from limited semantic information in the prompt representation or insufficient supervision for learning high-dimensional densities.
@@
Combining pre- and post-generation signals. SSD exposes complementary reliability signals from a single predicted density: pre-generation dispersion forecasts risk, while post-generation likelihood verifies prompt–answer compatibility. A natural next step is to integrate these signals into unified decision policies for LLM-based systems, for example abstaining or retrieving evidence under high dispersion, and accepting or rejecting candidate actions based on likelihood thresholds.
Diffusion models. Although our experiments focus on autoregressive models, SSD may be particularly well suited to diffusion-based language models [sahoodiffusion2024]. Since diffusion models do not rely on token-wise autoregressive decoding, the mapping from prompt representation to final output distribution may be more direct, potentially reducing the gap between the model’s internal belief state and the semantic distribution induced by sampling.
Domain-specific uncertainty estimation. As described in Section 2.3, SSD extends beyond language to sequence models whose outputs admit meaningful latent representations. We see particular potential utility for this technique in electronic health record (EHR) foundation models, which tokenize longitudinal patient histories and generate future clinical event sequences [renc2024zero]. These models typically predict discrete events, but clinical utility often depends on a higher-level characterization of the patient’s future state, such as overall disease burden or risk profile [steinberg2024motor, zeng2025trajsurvlearningcontinuouslatent]. In this setting, the embedding model could correspond to a transformer-based EHR encoder such as BEHRT [li2020behrt], summarizing a predicted trajectory into a clinically informative latent health representation. Sampling from $q_{\phi}(\mathbf{z}\mid\mathbf{h})$ could then yield low-latency forecasts of plausible future patient states, while the entropy of the density would quantify uncertainty over long-term outcomes. Similar structure arises in time-series foundation models [faw2025incontext], where future trajectories may be summarized by latent representations capturing regime, trend, or seasonal structure [wang2022learning].
@@ 6 Conclusion
We introduced Semantic Self-Distillation (SSD), a framework for distilling a language model’s sampled semantic answer distribution into a lightweight, prompt-conditioned predictive density. Rather than reducing uncertainty to a scalar score, SSD models the geometry of the semantic output space itself, enabling analytic dispersion estimates and likelihood-based evaluation in a single forward pass.
Empirically, SSD provides competitive hallucination prediction performance relative to sampling-based semantic dispersion baselines across several model families, while avoiding the inference-time overhead of generating multiple completions. Crucially, modelling the full density enables capabilities unavailable to scalar probes, including answer-level reliability scoring and distribution-based selection.
More broadly, SSD demonstrates that uncertainty in complex generative models can be distilled as an explicit density over latent outcome representations. By shifting computational cost from inference to training, SSD provides a practical mechanism for integrating principled, distribution-level uncertainty into latency-critical language systems and other sequence domains.
Acknowledgements.
@@ Appendix A Derivations
Appendix B Further Analyses
In this section we provide additional analyses supporting observations made in the main text. Unless otherwise specified, results are reported on the TriviaQA dataset, where SSD exhibited the strongest performance variation across base models.
B.1 Distillation Fidelity vs. Detection Performance
Figure A1 plots hallucination detection AUROC against distillation fidelity across base models and student hyperparameter settings. We observe a strong positive relationship, with a cross-model Spearman correlation of $\rho_{\text{meta}}=0.65$, indicating that detection performance is limited by the student’s ability to learn the prompt-to-dispersion mapping. The failure cases, notably Qwen3-4B, exhibit near-zero $\rho_{\text{fidelity}}$, suggesting that the relevant uncertainty signal is not accessible from the chosen prompt representation.
Figure A1: Distillation fidelity drives detection performance. The X-axis shows distillation fidelity ($\rho_{\text{fidelity}}$): the Spearman correlation between the student’s predicted entropy and the teacher dispersion across TriviaQA test prompts. Multiple points per model correspond to different student hyperparameter settings. $\rho_{\text{meta}}$ is the Spearman correlation between the X and Y variables; it shows that models where the student fails to learn the distribution (low $\rho_{\text{fidelity}}$) result in poor detection.
B.2 Semantic Consensus from the Predicted Mixture
In Section 2.2, we noted that modelling an explicit predictive density enables latent-space sampling and computation of mixture statistics such as the mean. In language models, the mixture mean can serve as a proxy for semantic consensus. Whereas autoregressive decoding produces a single trajectory through output space, the mixture mean aggregates across the model’s sampled distribution in embedding space. We evaluate whether this analytic mean more faithfully approximates the empirical semantic centroid of sampled answers than a single low-temperature decoding.
We treat the empirical mean of $S=32$ stochastic samples in embedding space as a reference semantic centroid, then measure the mean squared distance (MSD) in Euclidean space between this centroid and both (i) the model’s low-temperature output and (ii) the analytical mean of the student mixture, $\bm{\mu}_{ssd}=\sum_{k}\pi_{k}\bm{\mu}_{k}$. Lower MSD indicates better recovery of the model’s aggregate semantic belief. MSD is computed as $\mathbb{E}_{i}\left[\|\hat{\mathbf{z}}_{i}-\bm{\mu}^{\text{true}}_{i}\|_{2}^{2}\right]$ over examples $i$ in the subset.
We report results for the best-performing student configuration (Ministral 8B), reflecting the strong architectural dependence of distillation fidelity. As shown in Table A1, the SSD mixture mean provides a closer approximation to the semantic centroid than a single default decoding pass. This effect is most pronounced for incorrect answers, where SSD improves centroid recovery in over 76% of test cases. This aligns with prior work showing that hallucinated outputs tend to deviate from semantic consensus [lin2024generating, phillips2025geometricuncertaintydetectingcorrecting]. Together with the OOD results in Table 3, this result suggests that SSD exposes a coherent semantic belief state from which both prior (entropy) and posterior (likelihood and distance-to-consensus) reliability signals can be derived.
Table A1: Semantic consensus approximation with Ministral 8B. We report the mean squared distance (MSD) of the default answer embedding and the SSD predicted mean to the true semantic centroid (average of $S=32$ samples). Values denote mean MSD, with the bootstrap standard deviation expressed as a percentage of the mean shown as a subscript.
All 0.0153.2 0.0081.9 +45.5 58.3
Correct 0.0134.4 0.0092.3 +32.3 48.8
Incorrect 0.0184.2 0.0073.4 +63.0 76.5
Appendix C Ablations
We analyze two design axes underlying SSD: the dimensionality of the semantic embedding space and the capacity of the student mixture model. All ablations are conducted using the TriviaQA dataset.
C.1 PCA Dimension
@@
This introduces a trade-off between semantic resolution and distillation fidelity: higher dimensions preserve more semantic detail (raising the teacher dispersion baseline), but make the prompt-to-density regression task harder for the student to learn with fixed data. As shown in Figure A2 and Table A2, while the detection performance of teacher dispersion improves monotonically with dimensionality, the student’s performance (SSD) peaks at $d_{pca}=16$. Beyond this point, the student fails to effectively distil the increasingly sparse density, leading to a drop in detection accuracy. For the OOD experiments we find PCA dimensions of 16 and 32 perform similarly, followed by a drop-off at higher dimensions. We fix $d_{pca}=16$ for all main experiments bar the OOD task, where we set $d_{pca}=32$.
@@
Table A2: Effect of PCA dimensionality on Hallucination prediction (SSD) and OOD Verification. Scores are averaged across the best student configuration for each model.
@@
Figure A2: As PCA dimension increases, the semantic resolution improves and the teacher dispersion baseline rises. The distillation task however becomes harder, causing SSD performance to fall.
C.2 Student Capacity
@@
Table A4 reveals that the optimal hyperparameters vary by base model. Aggregating results across all models (Table A3) reveals capacity saturation, where increasing the network width or depth yields negligible gains in mean AUROC. This saturation indicates that performance is not limited by student expressivity, but by the information available in the prompt representation used for distillation. The student appears to extract most of the accessible uncertainty signal from this representation; further improvements would require richer inputs (e.g., attention-pooled representations) rather than larger students. Although we observe optimal performance for $K=10$ mixture components, we opt not to increase this further given our set number of $S=32$ samples; modelling more than ten unique semantic concepts in this setting is unrealistic and would likely damage the generalization ability of the trained student.
Table A3: Marginal impact of student capacity (Fixed PCA=16). Scores represent the mean AUROC across all runs with the specified parameter; standard deviations indicate variability across different base models and remaining hyperparameter configurations. We observe negligible performance gains from increasing complexity, indicating information saturation in the prompt representation.
@@
Figure A3: Student capacity ablation on the TriviaQA dataset, averaged across all seven models investigated.
Appendix D Model Configuration and Selection
Here we summarize model characteristics and specific configuration choices used throughout our experiments. Table A4 reports base model accuracy on TriviaQA, the optimal probe layers identified for PCP and SEP, and the student architecture that achieved the highest SSD AUROC for each model.
Table A4: Model-specific configuration details on TriviaQA. ‘Accuracy’ denotes base model QA accuracy. ‘Total’ gives the number of transformer layers, while ‘PCP’ and ‘SEP’ report the selected probe layers. ‘Best SSD Config’ lists the student architecture (HH=hidden size, DD=depth, KK=mixture components) achieving the highest SSD AUROC.
Layers Best SSD Config
Model Accuracy Total PCP SEP H D K AUROC
Qwen3-8B 0.620 36 32 24 128 6 10 0.677
Qwen3-4B 0.565 36 18 20 256 2 10 0.699
Llama 3.1 8B 0.700 32 20 16 256 4 2 0.687
Llama 3.2 3B 0.585 28 15 19 512 3 5 0.737
Ministral 8B 0.656 36 33 34 1024 6 1 0.748
SmolLM3 3B 0.567 36 20 19 512 6 5 0.710
Gemma 3 4B 0.532 34 22 19 512 2 10 0.700
D.1 Layer Selection Protocol
To select the student input layer $\ell^{*}$, we follow kossen2024semanticentropyprobesrobust and train linear probes on all layers to predict binarized Semantic Entropy. We choose the layer that maximizes validation accuracy on this proxy task. This heuristic assumes that layers most predictive of scalar dispersion also contain the richest signal for learning the full semantic density. In practice, the selected layers tend to lie in deeper parts of the network (Table A4).
Appendix E Computational Complexity Analysis
Sampling-based Methods (SE, TD).
Computing Semantic Entropy [farquhar_detecting_2024] requires $S$ stochastic sequences per prompt at inference time (typically $S\in[5,20]$). The cost is dominated by the autoregressive decoding steps:
where $\mathcal{O}_{\text{NLI}}$ represents the quadratic cost of comparing all sample pairs using an entailment model. This linear scaling with $S$ introduces latency that is unacceptable in many applications, for example in interactive agentic systems.
Scalar Probes (PCP, SEP).
Semantic Self-Distillation (SSD).
Appendix F Full Hallucination Prediction Results
We report full per-model hallucination detection results corresponding to the summaries in Tables 2 and 1. These tables provide AUROC and AUPRC metrics for all methods across datasets.
Table A5: Hallucination Prediction AUROC. Values report mean AUROC (%) over 1,000 bootstrap resamples of the test set, with the bootstrap standard deviation shown as a subscript. We mark the best method per dataset and model in bold.
TriviaQA MMLU
Model PCP SEP SE TD SSD PCP SEP SE TD SSD
Qwen3 8B 76.91.5 76.21.5 81.61.4 68.31.6 67.71.8 69.81.6 59.51.8 63.71.6 63.51.7 57.11.8
Qwen3 4B 80.21.4 78.31.4 78.61.4 73.21.6 69.91.6 67.91.7 64.31.8 60.51.6 63.11.8 64.91.7
Llama 3.1 8B 77.01.7 74.41.7 85.11.3 69.61.7 68.71.9 71.51.6 64.91.7 70.61.6 69.01.6 63.81.7
Llama 3.2 3B 76.71.5 75.71.6 80.61.4 66.61.7 73.71.6 66.01.9 64.41.8 70.11.7 67.11.8 63.81.7
Ministral 8B 74.01.7 77.11.6 80.51.5 74.31.7 74.91.6 67.01.7 62.81.8 69.31.7 65.11.8 63.61.8
SmolLM3 3B 79.61.4 78.71.4 83.11.2 69.91.6 71.11.6 67.01.7 66.31.8 68.81.7 62.31.8 62.01.8
Gemma 3 4B 75.21.6 70.61.6 73.11.5 67.61.6 70.01.7 66.21.8 57.81.8 61.31.6 62.81.7 60.71.8
Table A6: Hallucination Prediction AUPRC. Values report mean AUPRC (%) over 1,000 bootstrap resamples of the test set, with the bootstrap standard deviation shown as a subscript. We mark the best method per dataset and model in bold.
TriviaQA MMLU
Model PCP SEP SE TD SSD PCP SEP SE TD SSD
Qwen3 8B 64.82.6 64.92.7 75.02.1 52.42.6 56.92.8 66.92.3 56.32.3 60.32.2 61.92.3 55.92.3
Qwen3 4B 74.32.4 70.82.5 73.42.2 59.72.4 61.92.5 60.12.5 57.62.5 53.42.3 58.52.4 59.42.5
Llama 3.1 8B 59.73.1 54.03.0 73.42.7 44.42.7 51.92.9 72.72.1 62.62.3 70.52.1 70.92.1 62.72.4
Llama 3.2 3B 68.92.3 65.42.6 74.12.2 55.82.6 65.52.5 76.41.9 74.52.0 78.41.8 74.02.0 74.51.9
Ministral 8B 59.23.0 62.92.8 69.52.8 57.92.8 59.92.8 72.92.2 70.72.1 76.01.9 71.42.1 71.72.1
SmolLM3 3B 75.02.1 73.12.2 78.81.9 57.12.5 62.42.5 67.72.3 67.62.2 70.52.2 64.62.2 63.02.3
Gemma 3 4B 69.22.5 64.72.5 69.22.2 63.12.3 66.12.5 73.52.1 63.82.1 67.61.9 71.12.0 69.22.1